# InstanceGaussian: Appearance-Semantic Joint Gaussian Representation for 3D Instance-Level Perception

### v1

@@ 1 Introduction
footnotetext: *Corresponding author
@@
To address the aforementioned challenges, we propose InstanceGaussian, a method capable of jointly learning object appearance and semantic features while adaptively aggregating objects. Our contributions are summarized as follows:
@@
    Progressive Appearance-Semantic Joint Training Strategy: We propose a progressive training strategy that optimizes the joint representation of appearance and semantics in a step-by-step manner. This ensures their consistency throughout the training process, thereby enhancing the model’s stability and improving the accuracy of downstream tasks such as 3D point segmentation and 2D image segmentation.
@@ 3.1 Preliminary
3D Gaussian Splatting  [17]. Given a set of $n$ 3D Gaussians $\mathcal{G}$, each with center $\boldsymbol{\mu}$, color $\boldsymbol{c}$, opacity $\boldsymbol{\sigma}$, rotation $\boldsymbol{R}$, and scale $\boldsymbol{S}$, they are rendered as an image $\boldsymbol{I}_{ren}$ at a specific viewpoint via differentiable rasterization $\mathcal{D}$. The reconstruction loss $\mathcal{L}_{rgb}$ is then calculated by comparing $\boldsymbol{I}_{ren}$ with the GT image $\boldsymbol{I}_{gt}$. This process can be simplified as shown below, where blue variables are the optimizable parameters.
@@ 3.2 Appearance-semantic joint Gaussian representation
(1) Appearance-semantic balanced representation. For the same object or region, the texture details may vary while the semantics remain consistent. Consequently, many Gaussians with diverse appearance attributes are required to represent the texture details of the region. However, these Gaussians need only an identical semantic attribute to express the region’s semantics adequately. In other words, the same number of Gaussians may under-represent appearance but over-represent semantics. We refer to this phenomenon as the imbalance in the representation of appearance and semantics by Gaussians. To this end, we propose a Semantic-Scaffold-GS representation to balance appearance and semantics:
@@
The blue parts are the optimizable parameters; see Sec. 3.1 for variable definitions. Similar to Scaffold-GS, we derive $n$ neural Gaussian points $\mathcal{G}$ from $n^{\prime}$ anchor point $\mathcal{A}$ ($n=5n^{\prime}$ in our implementation). Specifically, each group of 5 child Gaussians with different appearance attributes is decoded from the appearance embedding $\boldsymbol{e}$ of their common parent anchor point using MLPs. However, at the semantic level, these 5 child Gaussians share the same semantic feature $\boldsymbol{f}$ as their parent anchor point. On the one hand, this representation strengthens the learning of appearance, especially in scenes represented by sparse points (e.g., ScanNet and SLAM-reconstructed scenes); on the other hand, sharing features among neighboring points significantly enhances the efficiency of semantic learning.
@@
(2) Appearance-semantic consistency learning. Without semantic constraints, i.e., pure appearance reconstruction, a Gaussian can represent across objects/regions. For instance, at boundaries, a single Gaussian may simultaneously represent both the object and the background. Recent language-embedded Gaussian Splatting works [26, 35] follow a paradigm of pre-training appearance reconstruction, then freezing the appearance properties, and subsequently learning semantic attributes. This tends to cause inconsistencies between appearance and semantics, posing challenges for 3D point segmentation and 2D image segmentation. Therefore, we propose a progressive appearance-semantic joint training strategy to ensure their consistency as shown in Fig. 2: during the first 0-10k steps, only the appearance attributes are trained, then appearance and semantics are trained independently in the next 10k-20k steps, and finally, appearance and semantics are jointly trained in the last 20k-30k steps. Here, we employ the loss (Eq. (1) (2)) of OpenGaussian to train semantic features, but its inter-mask contrastive loss value is too large, causing instability of joint training. We overcome this problem through the following improvements:
@@
where $\mathds{1}_{\|\bar{M}_{i}-\bar{M}_{j}\|^{2}<\tau}$ takes values of 1 or 0, aiming to truncate the loss between two objects whose dissimilarity (quantified by the L2 distance) exceeds a threshold $\tau$ (set to 0.4 in our implementation). This improvement is intuitive because there is no need to enforce that the features of two different objects be infinitely distant from each other.
@@ 3.3 Bottom-up category-agnostic instantiation
We employ a bottom-up method for category-agnostic instantiation. Specifically, we first over-segment the scene into numerous sub-objects. Then, we construct a graph of these sub-objects and perform aggregation based on the graph connectivity, adaptively obtaining several independent and complete objects. Alg. 1 describes this process.
@@
We first sample $s$ seed points from all $n$ Gaussians through FPS and use them as the initial centers of $s$ clusters, where $s$ is far greater than the number of objects in the scene, with a default value of 1000. Then we start the clustering process, updating the instance feature of all Gaussians and the $s$ clusters. This process can be formulated as:
@@
4 Experiments
@@ 4.1 Category-Agnostic 3D Instance Segmentation
Settings. 1) Task: Category-agnostic 3D instance segmentation aims to segment independent 3D objects/instances in a given point cloud (Gaussians can naturally be treated as points) scene without considering their semantic categories, ensuring that each point ultimately belongs to only one instance. 2) Dataset and Metrics: Following OpenGaussian, we evaluate using the ScanNet [8] dataset with GT instance-level annotations. More specifically, we calculate the instance-level mIoU and mAcc metrics for the 10 scenes selected by OpenGaussian. 3) Baseline: We mainly compare with Gaussian-based methods OpenGaussian [35] and GaussianGrouping [39] with RGB images and SAM masks as inputs. Additionally, we provide the performance of the depth-assisted back-projection method MaskClustering [38] as a reference.
@@
4.2 Open-Vocabulary Query Point Cloud Understanding
@@
4.3 Open Vocabulary 3D Object Selection
and Rendering
Result. Quantitative results are shown in the Tab.  3, which demonstrates that even when rendering 3D Gaussians onto 2D images for evaluation. As previously analyzed, LangSplat and LEGaussian, due to their poor 3D perception capabilities, fail to accurately select 3D Gaussians relevant to the text query, resulting in low-quality rendering results. The performance advantage of our method stems from two aspects. First, the bottom-up adaptive instantiation ensures good discrimination between objects. Second, the joint representation and learning of appearance and semantics avoid the appearance-semantic inconsistency present in OpenGaussian. The qualitative results in Fig. 5 more intuitively demonstrate this point. Our method renders objects with clearer boundaries, unlike OpenGaussian, which due to its appearance-semantic ambiguity, requires some Gaussians to represent multiple objects/regions, especially evident in the boundary areas.
Figure 6: Visualization results of the appearance color and semantic features of the Gaussians ellipsoid in the 3D viewer demonstrate that our method achieves superior appearance-semantic consistency. (a)(c): Results of our progressive joint training. (b)(d): Without joint training.
@@ 4.4.1 Gaussian representation and training strategy
We conducted the ablation studies as shown in Tab. 4 to demonstrate the effectiveness of our proposed Semantic-Scaffold-GS joint representation and progressive joint training strategy. The experiments are evaluated using the semantic and instance segmentation tasks on the ScanNet dataset. The joint representation achieves a balance between appearance and semantics, allowing the same anchor point to express different appearances while sharing the same semantics. The progressive joint training avoids inconsistencies in appearance and semantics, enhances the distinguishability between objects, and thus leads to improved instance segmentation and semantic understanding performance.
Fig. 6 more intuitively demonstrates our avoidance of inconsistency. The first and second rows respectively visualize the appearance color and semantic features of the 3D Gaussian ellipsoid in the 3D viewer, where the ellipse represents the boundary of the Gaussian. Fig. 6 (a)(c) shows the results of our joint training, which achieved clearer and more distinctive expressions in both appearance and semantics compared to Fig. 6 (b)(d). This improvement not only brings benefits to the understanding of the 3D level (i.e., the semantic level, as proven by our quantitative ablation studies), but also enhances the quality of object rendering (i.e., the appearance level, as shown in Sec. 4.3).
@@
Table 4: Performance of instance segmentation and semantic segmentation with different representation and training strategies on Scannet. Evaluated by mIoU.
Case Condition Semantic Seg. Instance Seg.
feature voxelize
@@
Table 5: Performance of instance segmentation and semantic segmentation with different connectivity condition strategies. Evaluated by mIoU.
@@ 4.4.2 Aggregation Condition
In the process of aggregating over-segmented sub-objects based on a connectivity graph, we considered two connectivity conditions: the L2 distance of semantic features, and the proximity judgment after voxelization.
We conducted an ablation study on its configuration as shown in Tab. 5. In case #1, we used the color of the Gaussians as a feature and considered their positions for aggregation, which resulted in poor performance. In case #2, we only considered semantic features and achieved a significant performance improvement, demonstrating the effectiveness of our feature learning. In case #3, we took both conditions into account and achieved the best performance, which is an intuitive idea that in aggregation, we need to consider both semantic similarity and spatial proximity.
5 Conclusion
In this paper, we introduce an appearance-semantic compact Gaussian representation to address the imbalance between appearance and semantic features using Gaussians. On one hand, this representation enhances the learning of appearance; on the other hand, sharing features among neighboring points significantly boosts the efficiency of semantic learning. The bottom-up category-agnostic instantiation adaptively determines the number of categories, cleverly avoiding the issue of over-segmentation or under-segmentation. In category-agnostic segmentation tasks, we significantly outperform previous models, achieving state-of-the-art (SOTA) performance in open vocabulary understanding for 3D point-level tasks.
@@ Appendix A Implementation Details
For ScanNet [8] dataset, we freeze the point cloud coordinates and disable 3DGS [17] densification. For LeRF [18] dataset, we optimized the point cloud coordinates and enabled 3DGS densification. We stop 3DGS densification in 10k steps.
@@ A.2 Training time
We train each scene on a single 24G 3090 GPU (with actual memory usage around 5 to 10 GB). For the LeRF dataset, each scene takes around 200 images and trains for approximately 70 minutes. For the ScanNet dataset, each scene takes around 100-300 images, and trains for approximately 35 minutes. We selected level 3 to extract SAM [19] mask.
@@ A.3 ScanNet Dataset Setting
Appendix B More Visual Results
To demonstrate the effectiveness of our approach, we conducted additional experiments on ScanNet to prove our performance on category-agnostic 3D instance segmentation (Fig.7) and open-vocabulary query point cloud understanding (Fig.8). We also performed experiments on GraspNet [9] dataset (Fig.9), and the results indicate the generalization capabilities of our method.

### v2

@@ 1 Introduction
footnotetext: * Corresponding authorfootnotetext: This work is supported by Guangdong Provincial Key Laboratory of Ultra High Definition Immersive Media Technology(Grant No. 2024B1212010006)
@@
To address these challenges, we propose InstanceGaussian, a method capable of jointly learning object appearance and instance features while adaptively aggregating objects. Our contributions are summarized as follows:
@@
    Progressive Appearance-Semantic Joint Training Strategy: We propose a progressive training strategy that optimizes the joint representation of appearance and semantics* in a step-by-step manner. This ensures their consistency throughout the training process, thereby enhancing the model’s stability and improving the accuracy of downstream tasks such as 3D point segmentation and 2D image segmentation.
@@
footnotetext: * The term “semantic” is used here to distinguish it from “appearance”. In subsequent sections and code implementations, this concept will be explicitly referred to as “instance features”.
@@ 3.1 Preliminary
3D Gaussian Splatting  [18]. Given a set of $n$ 3D Gaussians $\mathcal{G}$, each with center $\boldsymbol{\mu}$, color $\boldsymbol{c}$, opacity $\boldsymbol{\sigma}$, rotation $\boldsymbol{R}$ and scale $\boldsymbol{S}$, they are rendered as an image $\boldsymbol{I}_{ren}$ at a specific viewpoint via differentiable rasterization $\mathcal{D}$. The reconstruction loss $\mathcal{L}_{rgb}$ is then calculated by comparing $\boldsymbol{I}_{ren}$ with the GT image $\boldsymbol{I}_{gt}$. This process can be simplified as shown below, where blue variables are optimizable parameters.
@@ 3.2 Appearance-semantic joint Gaussian Representation
(1) Appearance-semantic balanced representation. For the same object or region, the texture details may vary while the semantics remain consistent. Consequently, many Gaussians with diverse appearance attributes are required to represent the texture details of the region. These Gaussians require only a single semantic attribute to effectively capture the region’s semantics, meaning the same number of Gaussians may under-represent appearance while over-representing semantics. We refer to this phenomenon as the imbalance in the representation of appearance and semantics by Gaussians. To this end, we propose a Semantic-Scaffold-GS representation to balance appearance and semantics:
@@
The blue parts are the optimizable parameters; see Sec. 3.1 for variable definitions. Similar to Scaffold-GS, we derive $n$ neural Gaussian points $\mathcal{G}$ from $n^{\prime}$ anchor points $\mathcal{A}$ ($n=5n^{\prime}$ in our implementation). Specifically, each group of 5 child Gaussians with different appearance attributes is decoded from the appearance embedding $\boldsymbol{e}$ of their common parent anchor point using MLPs. However, at the semantic level, these 5 child Gaussians share the same instance feature $\boldsymbol{f}$ as their parent anchor point. On the one hand, this representation strengthens the learning of appearance, especially in scenes represented by sparse points (e.g., ScanNet and SLAM-reconstructed scenes); on the other hand, sharing features among neighboring points significantly enhances the efficiency of semantic learning. The process is described in the top row of Fig. 1.
@@
(2) Appearance-semantic consistency learning. Without semantic constraints, i.e., pure appearance reconstruction, a Gaussian can represent across objects/regions. For instance, at boundaries, a single Gaussian may simultaneously represent both the object and the background. Recent language-embedded Gaussian Splatting works [28, 37] follow a paradigm of pre-training appearance reconstruction, freezing the appearance properties, and subsequently learning semantic attributes. This tends to cause inconsistencies between appearance and semantics, posing challenges for 3D point segmentation and 2D image segmentation. Therefore, we propose a progressive appearance-semantic joint training strategy to ensure their consistency as shown in Fig. 2: during the 0-10k steps, only the appearance attributes are trained, then appearance and semantics are trained independently in the 10k-20k steps, and finally, appearance and semantics are jointly trained in the 20k-30k steps. Here, we employ the loss (Eq. (1) (2)) of OpenGaussian to train instance features, but its inter-mask contrastive loss value is too large, causing instability of joint training. We overcome this problem through the following improvements:
@@
where $\mathds{1}_{\|\bar{M}_{i}-\bar{M}_{j}\|^{2}<\tau}$ takes values of 1 or 0, aiming to truncate the loss between two objects whose dissimilarity (quantified by the L2 distance) exceeds a threshold $\tau$ (set to 0.4 in our implementation). This improvement is intuitive because there is no need to enforce that the features of two different objects are infinitely distant from each other.
@@ 3.3 Bottom-up Category-Agnostic Instantiation
We employ a bottom-up method for category-agnostic instantiation. Specifically, we first over-segment the scene into numerous sub-objects. Then, we construct a graph of these sub-objects and perform aggregation based on the graph connectivity, adaptively obtaining several independent and complete objects. The bottom row of Fig. 1 describes this process.
@@
We first sample $s$ seed points from all $n$ Gaussians through FPS and use them as the initial centers of $s$ clusters, where $s$ is much greater than the number of objects in the scene, with a default value of 1000. Then we start the clustering process, updating the instance feature of all Gaussians and the $s$ clusters. This process can be formulated as:
@@
4 Experiments*
footnotetext: * Please refer to the supplementary materials for additional 3D/2D visualization results and comparisons with more methods.
@@ 4.1 Category-Agnostic 3D Instance Segmentation
Settings. 1) Task: Category-agnostic 3D instance segmentation aims to segment independent 3D objects/instances in a given point cloud (Gaussians can naturally be treated as points) scene without considering their semantic categories, ensuring that each point ultimately belongs to only one instance. 2) Dataset and Metrics: Following OpenGaussian, we evaluate using the ScanNet [9] dataset with GT instance-level annotations. More specifically, we calculate the instance-level mIoU and mAcc metrics for the 10 scenes selected by OpenGaussian. 3) Baseline: We mainly compare with Gaussian-based methods OpenGaussian [37] and GaussianGrouping [41] with RGB images and SAM masks as input. Additionally, we provide the performance of the depth-assisted back-projection method MaskClustering [40] as a reference.
@@
4.2 Open-Vocabulary Query Point Cloud Understanding
@@
4.3 Open Vocabulary 3D Object Selection and
Rendering
Result. Quantitative results are displayed in the Tab. 3, which demonstrates that even when rendering 3D Gaussians onto 2D images for evaluation. As previously analyzed, LangSplat and LEGaussian, due to their poor 3D perception capabilities, fail to accurately select 3D Gaussians relevant to the text query, resulting in low-quality rendering results. The performance advantage of our method stems from two aspects. First, the bottom-up adaptive instantiation ensures good discrimination between objects. Second, the joint representation and learning of appearance and semantics avoid the appearance-semantic inconsistency present in OpenGaussian. The qualitative results in Fig. 5 more intuitively demonstrate this point. Our method renders objects with clearer boundaries, unlike OpenGaussian, which due to its appearance-semantic ambiguity, requires some Gaussians to represent multiple objects/regions, especially evident in the boundary areas.
Figure 6: Visualization results of the appearance color and instance features of the Gaussians ellipsoid in the 3D viewer demonstrate that our method achieves superior appearance-semantic consistency. (a)(c): Results of progressive joint training. (b)(d): Without joint training.
@@ 4.4.1 Gaussian Representation and Training Strategy
We conduct the ablation studies as shown in Tab. 4 to demonstrate the effectiveness of our Semantic-Scaffold-GS joint representation and progressive joint training strategy. The experiments are evaluated using the semantic and instance segmentation tasks on the ScanNet dataset. The joint representation achieves a balance between appearance and semantics, allowing the same anchor point to express different appearances while sharing the same semantics. The progressive joint training avoids inconsistencies in appearance and semantics, enhances the distinguishability between objects, and thus leads to improved instance segmentation and semantic understanding performance.
Fig. 6 more intuitively demonstrates our avoidance of inconsistency. The first and second rows respectively visualize the appearance color and instance features of the 3D Gaussian ellipsoid in the 3D viewer, where the ellipse represents the boundary of the Gaussian. Fig. 6 (a)(c) shows the results of our joint training, which achieve clearer and more distinctive expressions in both appearance and semantics compared to Fig. 6 (b)(d). This improvement not only brings benefits to the understanding of the 3D level (i.e., the semantic level, as proven by our quantitative ablation studies), but also enhances the quality of object rendering (i.e., the appearance level, as shown in Sec. 4.3).
@@
Table 4: Performance of instance and semantic segmentation with different representation and training strategies on Scannet. Evaluated by mIoU.
Case feature voxelize Semantic Seg. Instance Seg.
@@
Table 5: Performance of instance and semantic segmentation with different connectivity condition strategies. Evaluated by mIoU.
@@ 4.4.2 Aggregation Condition
In the process of aggregating over-segmented sub-objects based on a connectivity graph, we considered two connectivity conditions: the L2 distance of instance features, and the proximity judgment after voxelization.
We conduct an ablation study on its configuration as shown in Tab. 5. In case #1, we use the color of the Gaussians as a feature and consider their positions for aggregation, which results in poor performance. In case #2, we only consider instance features and achieve a significant performance improvement, demonstrating the effectiveness of our feature learning. In case #3, we take both conditions into account and achieve the best performance, which is an intuitive idea that in aggregation, we need to consider both semantic similarity and spatial proximity.
5 Conclusion and Limitation
In this paper, we introduce an appearance-semantic compact Gaussian representation to address the imbalance between appearance and instance features using Gaussians. On one hand, this representation enhances the learning of appearance; on the other hand, sharing features among neighboring points significantly boosts the efficiency of semantic learning. The bottom-up category-agnostic instantiation adaptively determines the number of categories, cleverly avoiding the issue of over-segmentation or under-segmentation. In category-agnostic segmentation tasks, we significantly outperform previous models, achieving state-of-the-art (SOTA) performance in open vocabulary understanding for 3D point-level tasks.
Despite achieving novel performance on multiple tasks, our method still has the following limitations: 1) Longer training time for slower rendering, since the appearance attributes need to be calculated through MLP. 2) Segmentation accuracy is still affected when most of the images in the scene are incorrectly segmented by SAM. Failure case analyses can be found in the supplementary materials. We hope those drawbacks can be addressed in future works.
@@ Appendix A Implementation Details
For ScanNet [9] dataset, we freeze the point cloud coordinates and disable 3DGS [18] densification. For the LeRF [19] dataset, we optimize the point cloud coordinates and enable 3DGS densification. We stop 3DGS densification in 10k steps.
@@ A.2 Training Time
We train each scene on a single 24G 3090 GPU (with actual memory usage around 5 to 10 GB). For the LeRF dataset, each scene takes around 200 images and trains for approximately 70 minutes. For the ScanNet dataset, each scene takes around 100-300 images and trains for approximately 30 minutes. Tab. 6 shows the comparison with the baseline method in training time. We selected level 3 to extract the SAM [20] mask.
Method Langslpat LEGaussian OpenGaussian Ours
Time (s) 1153 1011 1378 1847
Table 6: Comparation of training time in Scannet.
@@ A.3 ScanNet Dataset Setting
Appendix B Comparison with more Methods
We conducted comparisons with additional baselines in Tab. 7, demonstrating the superior performance of our method.
Semantic seg. Instance seg.
Method mIoU mAcc. mIoU mAcc.
GAGA[23] -unsupported- 16.76 31.49
SAGA[7] 9.44 16.23 34.16 70.14
Ours 40.66 54.01 50.27 80.22
Table 7: Comparison with more methods.
Appendix C Ablation Study
C.1 More Detailed Ablation Metrics
We conducted a detailed ablation study with mA50 and mA25 metric in Tab. 8 and Tab. 9. We chose joint representation and joint training as the best strategy for the segmentation topic, as mIoU is generally more reasonable than mAcc.
Joint Semantic seg Instance Seg
Rep. Train. mIoU mAcc mIoU mA50 mA25
30.71 44.22 47.4 44.51 84.42
33.15 44.45 49.57 52.82 82.17
40.66 54.01 50.27 52.57 80.22
Table 8: Detailed ablation of joint representation and training.
Condition Semantic seg Instance Seg
Feat. Vox. mIoU mAcc mIoU mA50 mA25
21.75 33.39 27.50 16.16 49.33
28.98 40.58 43.41 40.63 72.50
40.66 54.01 50.27 52.57 80.22
Table 9: Detailed ablation of aggregation condition.
C.2 Hyperparameters
We conducted ablation experiments of hyperparameters in Tab. 10, demonstrating the robustness of our instantiated algorithm. The hyperparameters reported in our paper were not carefully tuned, and we discovered even better performance when adjusting them.
sample number voxel size connectivity thredshold
hyperparameters 200$\sim$1000 0.05$\sim$0.5 0.06$\sim$0.18
semantic seg. mIoU 37.87$\pm$2.79 38.96$\pm$1.70 38.48$\pm$2.17
instance seg. mIoU 49.60$\pm$1.31 49.56$\pm$0.71 49.90$\pm$1.29
Table 10: Abliation of hyperparameters.
C.3 Randomness and Runtime of FPS
We randomly sampled different starting points for FPS in Tab. 11, and the results demonstrated the robustness of our method. We also evaluated runtime across different sample number ($K$) in Tab. 12. While the merging algorithm has a complexity of O($K^{2}$), $K$ shows limited impact on overall runtime in practice (up to 3.43%).
Sample times Semantic Seg. mIoU Instance Seg. mIoU
5 39.44$\pm$1.38 48.57$\pm$0.98
Table 11: Ablation experiment of randomness in FPS.
Variation Merging time(s) Total time(s) Percentage
200$\sim$1000 5$\sim$63 1772$\sim$1847 0.28$\sim$3.43%
Table 12: Ablation experiment of different $K$.
C.4 Aggregation Algorithm
We compare our aggregation algorithm against clustering features by HDBSCAN[5] in Tab. 13, demonstrating our superiority.
Method Semantic Seg. mIoU Instance Seg. mIoU
HDBSCAN 31.84 44.76
Ours 40.66 50.27
Table 13: Ablation of aggregation algorithm.
Appendix D More Visual Results
To demonstrate the effectiveness of our approach, we conducted additional experiments on ScanNet[9] to prove our performance on category-agnostic 3D instance segmentation (Fig. 7) and open-vocabulary query point cloud understanding (Fig. 8). Additionally, we provide more 2D instance segmentation results on LeRF[19] (Fig. 9). We also performed experiments on GraspNet [10] dataset (Fig. 10), and the results indicate the generalization capabilities of our method.